Computation of Variable Cost Variances
The following information pertains to the standard costs and actual activity for Tyler Company for September:

Standard cost per unit
Direct materials 4 units of material A × $2.00 per unit
1 unit of material B × $3.00 per unit
Direct labor 3 hours × $8.00 per hour
Activity for September
Materials purchased
Material A 3,400 units × $2.05 per unit
Material B 800 units × $3.10 per unit
Materials used
Material A 3,050 units
Material B 705 units
Direct labor used 2,050 hours × $8.30 per hour
Production output 700 units

There were no beginning direct materials inventories.

(a) Determine the materials price and quantity variances.

Material A Material B
Materials price variance Answer

Incorrect
Mark 0.00 out of 1.00

AnswerFU

Correct
Mark 1.00 out of 1.00

Answer

Incorrect
Mark 0.00 out of 1.00

AnswerFU

Correct
Mark 1.00 out of 1.00

Materials quantity variance Answer

Incorrect
Mark 0.00 out of 1.00

AnswerFU

Correct
Mark 1.00 out of 1.00

Answer

Incorrect
Mark 0.00 out of 1.00

AnswerFU

Correct
Mark 1.00 out of 1.00

(b) Determine the labor rate and efficiency variances.

Labor rate variance Answer

Incorrect
Mark 0.00 out of 1.00

AnswerFU

Correct
Mark 1.00 out of 1.00

Labor efficiency variance Answer

Incorrect
Mark 0.00 out of 1.00

AnswerFU

Correct
Mark 1.00 out of 1.00
Answer #1
Material A Material B
Material price Variance 170UF 80UF
Material Quantity Variance 500UF 15UF
Labour Rate Variance 615 UF
Labour Efficiency Variance 400F

Detail Working for your Easy Understanding

Material A
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($2/Unit - $2.05/Unit ) × 3400Unit = $170 U F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((4Units    X 700 Unit)-3050 Unit )X $2 = $500U F
Material B
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($3/Unit - $3.10/Unit ) × 800Unit = $80 U F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((1Units    X 700 Unit)-705 Unit )X $3 = $15 UF
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($8/hour - $8.30/hour) × 2050 Hours= $615 UF
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × standard Rate
= (3 Hour × 700Unit - 2050) × $8 = $400 F